# Vital Insight: Assisting Experts’ Context-Driven Sensemaking of Multi-modal Personal Tracking Data Using Visualization and Human-In-The-Loop LLM Agents

### v2

@@ 1. Introduction
In this paper, we present a user-centered design process to explore the complex sensemaking process of experts when interpreting multi-modal health sensing data and the opportunity to support this process. First, we conduct formative interviews with 12 experts [^1] We define “experts” as people who are experienced researchers in ubiquitous computing who have utilized passive sensing data for health-related applications. We will discuss in detail in Section 3.1 to answer RQ1: What are the challenges and needs when ’making sense’ of personal tracking data in current practices? (Study 1). Guided by our findings, we develop an initial prototype, Vital Insight version 1 (VI-1), that leveraged both visualization and Large Language Models to assist experts’ sensemaking. We then conduct an exploratory user testing session with 13 experts using VI-1 as a technology probe to perform several sensemaking tasks (Study 2). Based on the user testing and feedback on VI-1, we iterate on our prototype and designed Vital Insight version 2 (VI-2), mainly refining the LLM augmentation to include more human-in-the-loop guidance. To further evaluate the modified AI components, we conduct an A/B testing session with 20 experts to investigate whether the AI augmentation is indeed helpful in their sensemaking process (Study 3). Our findings from the two studies indicate that the experts found VI to be a valuable prototype for assisting them in conducting evidence-based sensemaking, giving high ratings in dimensions such as trustworthiness and helpfulness. Experts also found their sensemaking tasks more satisfying with AI components than without them.
@@
  • Uncover the sensemaking model of experts when dealing with complex personal tracking data
  • Design and develop a human-in-the-loop prototype to assist experts in understanding personal tracking data through visualization and LLM
@@ 2.2. Tools to assist the sensemaking of personal tracking data
Most of these tools mentioned above focus on better data representation as a means to help experts and users understand and reflect on the data. However, some tools recognize that simply presenting data is not enough for meaningful understanding, and leverage AI to provide automatic assistance. For example, machine learning models have been widely used to attach labels to sensing data and have proven successful in tasks such as activity recognition, stress detection and mental health (mishra2020evaluating, ; mishra2020continuous, ; 10.1145/3351274, ), sleep analysis (kristiansen2021machine, ) and more. However, most of these tools are limited to ground truth labels and can only perform basic recognition tasks, failing to assist in more complex levels of understanding of personal tracking data. While these labels serve as good recognition results, they fall short in supporting experts in their sensemaking process. With the emergence of large language models (LLMs), researchers have begun adopting their “common sense” knowledge to provide higher-level inferences on sensor data (cosentino2024towards, ; fang2024physiollm, ; ji2024hargpt, ; ouyang2024llmsense, ; fang2024physiollm, ). However, these tools are often designed not to assist experts but to compare their accuracy with experts’ results. As a result, due to issues like hallucinations, experts often struggle to trust and validate these results, making manual collaboration with such systems challenging (lu2023human, ; chiang2023can, ). We recognize that there is still a lack of tools that leverage both evidence and helpful inference to assist experts in the sensemaking process of personal tracking data.
@@ 3.2.2. Needs and Opportunities in assisting experts
Now, recognizing the iterative nature of the sensemaking process and the two distinct paths it can take, we must consider how we can support this process effectively. We have identified two primary needs based on these paths: direct representation of data and indirect inference. Experts require a direct representation of sensing data to manage the vast volumes of information, as sensor data often does not come in a human-readable format. Throughout our interviews, experts commonly relied on visualizations for this purpose. However, most visualizations lack systematic design, are overly generic, and are often created ad-hoc for specific tasks. Experts mentioned techniques like stacking modalities or creating correlation graphs, but these efforts are sporadic and have not been standardized even within the same research group. Beyond direct data representation, participants also expressed a need for assistance with automatic inference generation as potential guidance for their sensemaking. Even with effective visualizations, the transformation of data into actionable insights is still predominantly a manual process that the experts conduct. P9 highlighted a challenge,
“We just have so much data, and it’d be pretty overwhelming to just kinda put that all in there.” (P9)
In longitudinal studies that span months or even years, it becomes impractical for experts to examine every piece of data to get the desired granularity required for some applications like labels for machine learning (P12/13). As a result, assistance in quickly navigating to the desired information seemed crucial. Our participants noted a significant lack of tools for this purpose, as the sensemaking process is complex and often beyond the capabilities of current algorithms and tools. For instance, some experts who used machine learning in their studies mentioned that they only viewed benchmarks as “results to be compared” against ground truth data, rather than as useful assistance for them. However, they expressed interest in getting some extracted data insights that go beyond simple labels. Based on the insights from the interviews, we designed and developed our initial prototype.
@@ 4. Initial Prototype: Connect data with insights
In this section, we present the proposed design of our initial prototype, Vital Insight version 1 (VI-1), based on the four sensemaking scenarios and two primary goals. The design of VI-1 aimed to reflect the insights from the preliminary interviews as a prototype. VI-1 will be used as a technology probe to assist experts in different sensemaking tasks to generate high-level insights (hutchinson2003technology, ).
@@ 4.1. Data types and system infrastructure
Figure 1. Eight components in VI-1 powered by visualization and LLM: time selection, main visualization, user profile, user check-in (user conversation with chatbot), LLM-generated daily summary, LLM-generated question, LLM-anomaly detection, LLM-generated hourly summary
@@ 4.2. Interface design
The Time Selection panel (top left) allows users to select the start and end date/time to zoom in/out, and iteratively switch between a holistic view and details on demand, in line with best practices for visual information seeking through data visualization (shneiderman2003eyes, ; wootton2024charting, ). The User Profile panel on the top right consists of basic demographic details such as age and living situation, designed to provide experts with a holistic understanding of the participant’s lifestyle.
@@
The Main Visualization panel serves as the central component to represent the sensing data (Fig.  1 left). The visualization component features five categories of information gathered from the phone, smartwatch, and chatbot: location, health, phone usage, activity, and events. We used a small multiple-style time series plot arrangement to display each category of information along a unified timeline. Small multiples are a visualization technique where various attributes are plotted along a consistent x-axis scale, facilitating multifaceted exploration (bavoil2005vistrails, ), pattern recognition (lekschas2017hipiler, ), and general data exploration (van2013small, ). This method is particularly effective for tasks that involve direct visual comparisons of time series data, aiding in comparing, exploring, and analyzing trends (javed2010graphical, ; boyandin2012qualitative, ; bach2015small, ; meyer2010pathline, ; heer2009sizing, ). The Location in the background shows location labels, e.g., home and church. We generate this data by cross-referencing the raw GPS coordinates from the phone app with participants’ self-reported frequently visited locations and addresses. The system marks each location with a unique background color across all views with a label at the top, aiming to provide contextual information when experts examine the data. In Fig. 1, the plotted location labels are home (light green) and church (light orange). The Health view presents heart rate and respiration data from the smartwatch using scatter plots in red and blue. Two black trend-lines overlayed on the graph represent the participant’s average heart rate and respiration over a one-month period. Exclamation marks (!) next to the scatter plot indicate detected anomalies. We present more information about the anomaly detection module in section 4.3. The Phone Usage view includes Wi-Fi (connected/disconnected, purple line chart), phone battery levels (0-100, green line chart), and phone interaction (unlock/lock events, solid/dashed grey rectangle overlaid). The Activity view reports step counts and basic activity recognition. A bar chart presents the step counts, with the height of the bar representing the number of steps taken during the time indicated by the width. The activity view presents basic activities from iPhone’s standard activity recognition API, using different colored rectangles (‘stationary’ in gray, ’walking’ in green, ’automotive’ in purple), indicating the activities and their respective time spans. Finally, the Event view includes significant events throughout the day, including phone calls (green) and interactions with the Alexa chatbot (orange), and their respective durations.
@@
Daily Summary panel, provides a high-level overview of the participant’s day with three tabs: LLM Summary, User Check-in, and Questions. The LLM Summary tab presents an LLM-generated summary of the entire day in paragraphs and bullet points for easy reading, thus providing experts the ability to quickly understand a person’s day. The User Check-in tab displays the real conversation between the chatbot and the participant (orange ‘Chatbot’ event in visualization). The Questions tab presents a list of questions and missing information that the system identifies as potentially helpful for data interpretation. This information can serve as an inspiration for experts to reflect on potential actions, particularly when engaging with different stakeholders, such as speaking directly with participants. We present detailed rationale for the LLM-generated content in Sec. 4.3.
@@
The Anomaly panel provides insights about unusual data points. For each anomaly, the panel displays the specific time range, description, and a possible reason, all generated by the LLM. Users can select the specific anomaly time to view the detailed information. The Hourly Summary panel provides LLM-generated interpretations for each hour of data. Users can select a specific time (e.g. 4-5 AM) to view the corresponding interpretation. Similar to the daily summary, this panel offers an objective summary of the sensor data with minimal inference, a more creative insights derived from the data, and questions that the LLM identifies as useful for interpreting that hour’s data. This panel offers different levels of inferences on potential events in detail.
@@ 4.3. LLM Augmentation
Our initial prototype provides four main LLM functionalities: the hourly summary, daily summary, question, and anomaly, all developed based on the OpenAI GPT-4 model through zero-shot Natural Language String method (Fig. 2(kim2024health, ; gruver2024large, ; liu2023large, ). We provide detailed prompts in the Supplementary Materials.
There are two primary forms of data input: Sensing Data and Contextual Data. For Sensing Data, we process the data into semantic description. Discrete data, such as WiFi and phone battery levels, are directly converted into sentences with values and timestamps (e.g. “The battery level of the person’s phone is [BATTERY] at [TIME].”). For time-series data like heart rate, encoding each data point individually would overwhelm our system. Instead, since these data are typically sampled at a fixed rate, we group them and send them in an array format like “The person’s respiration from [TIME] to [TIME] (10-second interval) is [VALUE1, VALUE2, …].”, and sort this narrative in chronological order. Another data input is Contextual data which includes User Profile - demographic and routine information, and User Check-in - which we process conversation in a time(From [TIME] to [TIME]) and utterance ([ROLE]:[UTTERANCE]) format.
@@
After transferring raw data into semantic descriptions, the prototype generates LLM low-level summary, which processes the most granular level of data and provides an hourly summary. The prompts consist of the following components: Goal, Data Interpretation Guidance, Data, User Profile, User Check-in, Historical Summary from GPT, and a JSON Formatter. The Goal outlines the task’s objective and explains the purpose of each input component. Data Interpretation Guidance offers expert advice on analyzing the data. Data includes one hour of sensing data in the semantic format along with the User Profile and User Check-in. The Historical Summary from GPT retrieves outputs from the previous hour’s LLM-generated summary as additional contextual information. The JSON Formatter specifies the desired JSON output format. It outputs summary, inference, and questions as described in Sec. 4.2.
@@
The prototype then processes the low-level summary into LLM High-Level sensemaking. Recognizing the daily patterns in activity, it aggregates the outputs from 24 hours of the hourly summary and other contextual information, to generate a similar output containing a summary, inference, and questions, but on a daily basis. The prompt includes the Goal (“You are tasked with summarizing a user’s daily activities based on the provided information…”), User Profile, User Check-in, Historical Summary from GPT (24-hour output from the hourly function), and the JSON Formatter. The system generates and formats the results in a human-readable way, both as a paragraph and in bullet points for easier digestion. It also summarizes information like questions at a higher daily level over a 24-hour period.
@@
Other than summary from the hourly data, LLM anomaly detector identifies potential anomalies in different modalities directly using the semantic descriptions of each sensing modality. After identifying potential anomalies, it then provides explanations of these anomalies using contextual information like user check-in. We provide detailed prompts for all the LLM functions in Supplementary Materials. While we iteratively adjusted these prompts to optimize outputs, we provided all LLM-generated results as-is and unaltered during user testing.
@@
Past research has used similar methods to embed sensing data into narratives before sending to LLMs (ji2024hargpt, ; cosentino2024towards, ; yang2024drhouse, ; kim2024health, ; ouyang2024llmsense, ; fang2024physiollm, ). While not the main focus of this paper, our method specifically distinguishes between different types of sensing data and identifies appropriate ways to handle each type. We also developed a multi-level pipeline that integrates contextual information at each step of the inference process, generating different inferences at various levels of granularity. At the same time, we realized that embedding sensing data into LLMs is an active and rapidly growing research area. Therefore, we designed VI-1 to keep the backend LLM and the frontend display of information relatively separate to ensure easy integration of advanced embedding methods into VI in the future.

### v3

@@ 1. Introduction
In this paper, we present a user-centered design process to explore the complex sensemaking process of experts when interpreting multi-modal health sensing data and the opportunity to support this process. First, we conduct formative interviews with 12 experts [^1] We define “experts” as people who are experienced researchers in ubiquitous computing who have utilized passive sensing data for health-related applications. We will discuss in detail in Section 3.1 to answer RQ1: What are the challenges and needs when ‘making sense’ of personal tracking data in current practices? (Study 1). Guided by our findings, we develop an initial prototype, Vital Insight version 1 (VI-1), that leveraged both visualization and Large Language Models to assist experts’ sensemaking. We then conduct an exploratory user testing session with 13 experts using VI-1 as a technology probe to perform several sensemaking tasks (Study 2). Based on the user testing and feedback on VI-1, we iterate on our prototype and designed Vital Insight version 2 (VI-2), mainly refining the LLM augmentation to include more human-in-the-loop guidance. To further evaluate the modified AI components, we conduct an A/B testing session with 20 experts to investigate whether the AI augmentation is indeed helpful in their sensemaking process (Study 3). Our findings from the two studies indicate that the experts found VI to be a valuable prototype for assisting them in conducting evidence-based sensemaking, giving high ratings in dimensions such as trustworthiness and helpfulness. Experts also found their sensemaking tasks more satisfying with AI components than without them.
@@
  • Construct the sensemaking model of experts when dealing with complex personal tracking data
@@ 2.2. Tools to assist the sensemaking of personal tracking data
In addition to visualizing routine/activity-related personal tracking data, there is a substantial body of work in the visualization community focused on time-series and multi-modal data. Kehrer et al. highlight that multi-modal scientific data often involves different acquisition modalities, varying spatial/temporal resolutions, and diverse data types (kehrer2012visualization, ). These differences create challenges in visualization design, particularly in aligning data for effective fusion and comparative analysis. One common solution is small multiples, a technique where various attributes are plotted along a consistent x-axis scale, which supports multifaceted exploration (bavoil2005vistrails, ), pattern recognition (lekschas2017hipiler, ), and data exploration (van2013small, ). This approach is especially useful for tasks that require direct visual comparisons of time-series data, aiding in the comparison, exploration, and analysis of trends (javed2010graphical, ; boyandin2012qualitative, ; bach2015small, ; meyer2010pathline, ; heer2009sizing, ). Researchers have also investigated specific visualization methods for time-series data, such as line graphs for trends, bar or circle charts for cumulative or cyclic data (matthews1997worm, ; harris1999information, ), and axis-based techniques, along with parallel coordinates for multivariate data (10.1145/967900.968153, ; 970ebe970164479f81ed8be97a37f73e, ; 9362264, ). In addition to crafting the visual representation of data, past work has also proposed principles for the interaction and customization with visualizations. For instance, Schneiderman’s mantra “Overview first, zoom and filter, then details-on- demand” describes how data should be presented on screen so that it is most effective for users (shneiderman2003eyes, ). Other visual interaction methods proposed by past works include explore, reconfigure, select, encode, abstract, elaborate, filter, and connect. (yi2007toward, ; blascheck2018exploration, ; fang2020survey, ).
@@
Most of these tools mentioned above focus on better data representation as a means to help experts and users understand and reflect on the data. However, some tools recognize that simply presenting data is not enough for meaningful understanding, and leverage AI to provide automatic assistance. For example, machine learning models have been widely used to attach labels to sensing data and have proven successful in tasks such as activity recognition, stress detection and mental health (mishra2020evaluating, ; mishra2020continuous, ; 10.1145/3351274, ), sleep analysis (kristiansen2021machine, ) and more. However, most of these tools are limited to ground truth labels and can only perform basic recognition tasks, failing to assist in more complex levels of understanding of personal tracking data. While these labels serve as good recognition results, they fall short in supporting experts in their sensemaking process. With the emergence of large language models (LLMs), researchers have begun adopting their “common sense” knowledge to provide higher-level inferences on sensor data (cosentino2024towards, ; fang2024physiollm, ; ji2024hargpt, ; ouyang2024llmsense, ; kim2024health, ). For instance, HARGPT uses LLMs to identify activities from raw IMU data (ji2024hargpt, ); Health-LLM evaluated zero/few-shot learning and fine-tuning on sensor data (kim2024health, ); and LLMSense and PhysioLLM use text-formatted sensor data and prompts to derive high-level inferences from sensor traces (ouyang2024llmsense, ; fang2024physiollm, ). However, these tools are often designed not to assist experts but to compare their accuracy with experts’ results. Due to issues like hallucinations, experts still struggle to trust and validate these results, making manual collaboration with such systems challenging (lu2023human, ; chiang2023can, ). Some recent works have begun to explore the use of AI as an aid to human sensemaking in fields such as genomics (mastrianni2024aienhancedsensemakingexploringdesign, ), social relationships (10.1145/3411764.3445290, ), annotation (10.1145/3411763.3451798, ), stroke diagnosis (jussupow2022radiologists, ), supply chain survival (sharma2025artificial, ), and more. Wenskovitch et al. further categorize AI’s roles in data sensemaking into four types: explorer, investigator, teacher, and judge, emphasizing the need for mutually intelligible, bidirectional communication (9619901, ). With the increasing focus on using AI as an assistance in the human sensemaking process, we recognize that there is still a lack of tools that leverage both evidence and helpful inferences to support experts in interpreting personal tracking data.
@@ 3.2.2. Needs and Opportunities in assisting experts
Now, recognizing the iterative nature of the sensemaking process and the two distinct paths it can take, we must consider how we can support this process effectively. We have identified two primary needs based on these paths: direct representation of data and indirect inference. Experts require a direct representation of sensing data to manage the vast volumes of information, as sensor data often does not come in a human-readable format. Throughout our interviews, experts commonly relied on visualizations for this purpose. However, most visualizations lack systematic design, are overly generic, and are often created ad-hoc for specific tasks. Experts mentioned techniques like stacking modalities or creating correlation graphs, but these efforts are sporadic and have not been standardized even within the same research group. Beyond direct data representation, participants also expressed a need for assistance with automatic inference generation as potential guidance for their sensemaking. Even with effective visualizations, the transformation of data into actionable insights is still predominantly a manual process that the experts conduct, which is overwhelming when dealing with so much data (P9). In longitudinal studies that span months or even years, it becomes impractical for experts to examine every piece of data to get the desired granularity required for some applications like labels for machine learning (P12/13). As a result, assistance in quickly navigating to the desired information seemed crucial. Our participants noted a significant lack of tools for this purpose, as the sensemaking process is complex and often beyond the capabilities of current algorithms and tools. For instance, some experts who used machine learning in their studies mentioned that they only viewed benchmarks as “results to be compared” against ground truth data, rather than as useful assistance for them. However, they expressed interest in getting some extracted data insights that go beyond simple labels. Based on these insights, we designed and developed our initial prototype.
@@ 4. Initial Prototype: Connect data with insights
In this section, we present the proposed design of our initial prototype, Vital Insight version 1 (VI-1), based on the four sensemaking scenarios and two primary goals. The design of VI-1 aimed to reflect the insights from the preliminary interviews as a prototype and underwent several iterations through casual group discussions with experts. VI-1 will be used as a technology probe to assist experts in different sensemaking tasks to generate high-level insights (hutchinson2003technology, ).
@@ 4.1. Data types and system infrastructure
Figure 1. Eight components in VI-1 powered by visualization and LLM: time selection, main visualization, user profile, user check-in (user conversation with chatbot), LLM daily summary, LLM question, LLM anomaly detection, LLM hourly summary
@@ 4.2. Interface design
The Time Selection panel (top left) allows users to select the start and end date/time to zoom in/out, and iteratively switch between a holistic view and details on demand to facilitate customization, in line with best practices for visual information seeking through data visualization (shneiderman2003eyes, ; wootton2024charting, ). The User Profile panel on the top right consists of basic demographic details such as age and living situation, designed to provide experts with a holistic understanding of the participant’s lifestyle (Scenario 1).
@@
The Main Visualization panel serves as the central component to represent the sensing data (Fig.  1 left). The visualization component features five categories of information gathered from the phone, smartwatch, and chatbot: location, health, phone usage, activity, and events. We used a small multiple-style time series plot arrangement to display each category of information along a unified timeline. The design also aligns with current practices, where experts manually stack different modalities of data, as described in the interview (Section 3.2.2). The Location in the background shows location labels, e.g., home and church. We generate this data by cross-referencing the raw GPS coordinates from the phone app with participants’ self-reported frequently visited locations and addresses. The system marks each location with a unique background color across all views with a label at the top, aiming to provide contextual information when experts examine the data. In Fig. 1, the plotted location labels are home (light green) and church (light orange). The Health view presents heart rate and respiration data from the smartwatch using scatter plots in red and blue. Two black trend-lines overlayed on the graph represent the participant’s average heart rate and respiration over a one-month period. Exclamation marks (!) next to the scatter plot indicate detected anomalies. We present more information about the anomaly detection module in section 4.3. The Phone Usage view includes Wi-Fi (connected/disconnected, purple line chart), phone battery levels (0-100, green line chart), and phone interaction (unlock/lock events, solid/dashed grey rectangle overlaid). The Activity view reports step counts and basic activity recognition. A bar chart presents the step counts, with the height of the bar representing the number of steps taken during the time indicated by the width. The activity view presents basic activities from iPhone’s standard activity recognition API, using different colored rectangles (‘stationary’ in gray, ’walking’ in green, ’automotive’ in purple), indicating the activities and their respective time spans. Finally, the Event view includes significant events throughout the day, including phone calls (green) and interactions with the Alexa chatbot (orange), and their respective durations.
@@
Daily Summary panel, provides a high-level overview of the participant’s day with three tabs: LLM Summary, User Check-in, and Questions. The LLM Summary tab presents an LLM-generated summary of the entire day in paragraphs and bullet points for easy reading, thus providing experts the ability to quickly understand a person’s day. We selected LLM to generate the summary due to its extensive capability to customize text-format summaries and explanations, offering different levels of detail and explainability for experts. The User Check-in tab displays the real conversation between the chatbot and the participant (orange ‘Chatbot’ event in visualization). The Questions tab presents a list of questions and missing information that the system identifies as potentially helpful for data interpretation. This information can serve as an inspiration for experts to reflect on potential actions, particularly when engaging with different stakeholders, such as speaking directly with participants (Scenario 4). We present detailed rationale for the LLM-generated content in Sec. 4.3.
@@
The Anomaly panel provides insights about anomalies and outliers, typically defined as “an observation that deviates significantly from other observations, raising suspicions that it was generated by a different mechanism,” and is crucial in analyzing time-series data such as heart rate (hawkins1980identification, ). In previous interviews, experts noted that anomalies are worth paying attention to (Scenario 3). For each anomaly, the panel displays the specific time range, description, and a possible reason, all generated by the LLM. Users can select the specific anomaly time to view the detailed information. The Hourly Summary panel provides LLM-generated interpretations for each hour of data. Users can select a specific time (e.g. 4-5 AM) to view the corresponding interpretation. Similar to the daily summary, this panel offers an objective summary of the sensor data with minimal inference, a more creative insights derived from the data, and questions that the LLM identifies as useful for interpreting that hour’s data. This panel offers different levels of inferences on potential events in detail (Scenario 2).
@@ 4.3. LLM Augmentation
Our initial prototype provides four main LLM functionalities: the hourly summary, daily summary, question, and anomaly, all developed based on the OpenAI GPT-4o-mini model through zero-shot Natural Language String method using static prompts (Fig. 2(kim2024health, ; gruver2024large, ; liu2023large, ). We provide detailed prompts in the Supplementary Materials.
There are two primary forms of data input: Sensing Data and Contextual Data. For Sensing Data, we process the data into semantic description before sending it to the model. Discrete data, such as WiFi and phone battery levels, are directly converted into sentences with values and timestamps (e.g. “The battery level of the person’s phone is [BATTERY] at [TIME].”). For time-series data like heart rate, encoding each data point individually would overwhelm our system. Instead, since these data are typically sampled at a fixed rate, we group them and send them in an array format like “The person’s respiration from [TIME] to [TIME] (10-second interval) is [VALUE1, VALUE2, …].”, and sort this narrative in chronological order. Another data input is Contextual data which includes User Profile - demographic and routine information, and User Check-in - which we process conversation in a time(From [TIME] to [TIME]) and utterance ([ROLE]:[UTTERANCE]) format. We do not specifically handle missing or noisy data and transfer each data point as-is.
@@
After transferring raw data into semantic descriptions, we send these descriptions to the model, which generates LLM low-level summary, processing the most granular level of data and providing an hourly summary. The prompts consist of the following components: Goal, Data Interpretation Guidance, Data, User Profile, User Check-in, Historical Summary from GPT, and a JSON Formatter. The Goal outlines the task’s objective and explains the purpose of each input component. Data Interpretation Guidance offers expert advice on analyzing the data. Data includes one hour of sensing data in the semantic format along with the User Profile and User Check-in. The Historical Summary from GPT retrieves outputs from the previous hour’s LLM-generated summary as additional contextual information. The JSON Formatter specifies the desired JSON output format. It outputs summary, inference, and questions as described in Sec. 4.2.
@@
The prototype then processes the low-level summary into LLM High-Level sensemaking. Recognizing the daily patterns in activity, it aggregates the outputs from 24 hours of the hourly summary and other contextual information, to generate a similar output containing a summary, inference, and questions, but on a daily basis. The prompt includes the Goal (“You are tasked with summarizing a user’s daily activities based on the provided information…”), User Profile, User Check-in, Historical Summary from GPT (24-hour output from the hourly function), and the JSON Formatter. The system generates and formats the results in a human-readable way, both as a paragraph and in bullet points for easier digestion. It also summarizes information like questions at a higher daily level over a 24-hour period. Before sending the summary to the dashboard, we perform several programmatic checks and rerun the process if the output does not meet the standard (e.g., output tokens¡10).
@@
Other than summary from the hourly data, LLM anomaly detector identifies potential anomalies in different modalities directly using the semantic descriptions of each sensing modality. While we iteratively adjusted these prompts to optimize outputs, we provided all LLM-generated results as-is and unaltered during user testing. If the inputs exceed token limits, we split them into chunks before sending them to the model; however, after this multi-level procedure, it’s very unlikely that the input tokens will exceed the context window of the model (128k).
@@
Past research has used similar methods to embed sensing data into narratives before sending to LLMs (ji2024hargpt, ; cosentino2024towards, ; yang2024drhouse, ; kim2024health, ; ouyang2024llmsense, ; fang2024physiollm, ). While not the main focus of this paper, our method specifically distinguishes between different types of sensing data and identifies appropriate ways to handle each type. We also developed a multi-level pipeline that integrates contextual information at each step of the inference process, generating different inferences at various levels of granularity. This multi-level pipeline also avoids sending an extensive amount of data to the model, which has been proved to be problematic and has led to inaccurate LLM outputs (levy2024same, ). At the same time, we realized that embedding sensing data into LLMs is an active and rapidly growing research area. Therefore, we designed VI-1 to keep the backend LLM and the frontend display of information relatively separate to ensure easy integration of advanced embedding methods into VI in the future.